OpenStudy (anonymous):

Solve the equation. Check your solution. "12x-28 = -63+7x" True or False?

OpenStudy (kl0723):

solve for x and then plug in the values back into the original eq

OpenStudy (kl0723):

12x-7x=-63+28 5x=-35 x=-35/5 x=-7

OpenStudy (kl0723):

Plug in the values: 12(-7)-28 = -63 +7(-7) left side = -112 right side = -112

OpenStudy (kl0723):

so I guess is true :P